OpenStudy (vera_ewing):

What is the graph of the system y = -2x + 3 and 2x + 4y = -8?

OpenStudy (vera_ewing):

@campbell_st I used an online grapher and none of the answer choices match the graph...

OpenStudy (campbell_st):

what are you asked to do..?

OpenStudy (vera_ewing):

What is the graph of the system y = -2x + 3 and 2x + 4y = -8?

OpenStudy (campbell_st):

ok... so the 1st line is in slop intercept form I'll put the 2nd line in slope intercept form 4y = -2x - 8 then y = -1/2x - 2 so you have 2 lines in slope intercept form the 1st line has a y- intercept of 3 the 2nd line has a y- intercept of -2 and both lines have negative slopes... use https://www.desmos.com/calculator to graph the lines.... and then compare them to your choices.... your answer is wrong as 1 line has a positive slope and the other a negative slope

OpenStudy (vera_ewing):

So then what would the answer be?

OpenStudy (campbell_st):

use the line to graph the lines... then compare it your choices...

OpenStudy (vera_ewing):

I know it's not B. I used an online grapher and it doesn't match with any of my choices at all!

OpenStudy (vera_ewing):

https://www.desmos.com/calculator/g4tx6bop27 It doesn't match any of them!

OpenStudy (campbell_st):

ok... that sounds odd here is my attempt using Desmos the equation of the lines are shown on the left.

OpenStudy (campbell_st):

perhaps you needed to zoom in so the scale on the axes is in 1 units

OpenStudy (vera_ewing):

That's really weird...so based on your graph, the answer is A?

OpenStudy (campbell_st):

yes... and my graph is the same as your graph, its just that I zoomed to have axes showing units of 1

OpenStudy (vera_ewing):

Oh ok thanks for explaing that!